# Curriculum learning based pre-training using Multi-Modal Contrastive Masked Autoencoders

### v1

@@ Abstract
In this paper, we propose a new pre-training method for image understanding tasks under Curriculum Learning (CL) paradigm which leverages RGB-D. The method utilizes Multi-Modal Contrastive Masked Autoencoder and Denoising techniques. Recent approaches either use masked autoencoding (e.g., MultiMAE [6]) or contrastive learning(e.g., Pri3D [45], or combine them in a single contrastive masked autoencoder architecture such as CMAE [47] and CAV-MAE [33]. However, none of the single contrastive masked autoencoder is applicable to RGB-D datasets. To improve the performance and efficacy of such methods, we propose a new pre-training strategy based on CL. Specifically, in the first stage, we pre-train the model using contrastive learning to learn cross-modal representations. In the second stage, we initialize the modality-specific encoders using the weights from the first stage and then pre-train the model using masked autoencoding and denoising/noise prediction used in diffusion models. Masked autoencoding focuses on reconstructing the missing patches in the input modality using local spatial correlations, while denoising learns high frequency components of the input data. Our approach is scalable, robust and suitable for pre-training with limited RGB-D datasets. Extensive experiments on multiple datasets such as ScanNet, NYUv2 and SUN RGB-D show the efficacy and superior performance of our approach. Specifically, we show an improvement of +1.0% mIoU against Mask3D on ScanNet semantic segmentation. We further demonstrate the effectiveness of our approach in low-data regime by evaluating it for semantic segmentation task against the state-of-the-art methods.
@@ 1 Introduction
Self-supervised learning (SSL) has paved the way for improving performance in various computer vision tasks such as object recognition [38, 87, 9], semantic segmentation [43, 29, 53] and depth estimation [35] by leveraging extensive unlabeled data and large models such as Vision Transformers (ViT) [26, 57]. There are two self-supervised frameworks commonly used to pre-train large vision models. First, contrastive learning [17, 89, 19, 34] pre-trains a model by maximizing the similarity between different augmentations of an image while minimizing the similarity to the other data samples. It essentially learns discriminative representations by enforcing invariance to data augmentations. Second family of SSL is Masked Image Modeling (MIM) which is inspired by the recent success of masking in the NLP domain [23]. The seminal work in MIM is Masked Autoencoders (MAE) [38] which masks out high number of patches in the input before passing them into a Transformer encoder. Then, a lightweight decoder is used to reconstruct the masked patches in the input. Such reconstruction task provides the model with understanding of the data beyond low-level image statistics. It has achieved state-of-art performance in a variety of image recognition tasks.
@@
Recently, Masked autoencoders have been extended to pre-train transformers using multi-modal data for tasks such as video-text retrieval [56, 83], image-text retrieval [68] and video-audio-text understanding [1]. However, utilizing MAEs for dense prediction tasks like semantic segmentation and depth estimation requires understanding beyond image level statistics. With the availability of RGB-D or Time-of-Flight (ToF) cameras, various approaches [6, 43, 31, 27] have been developed to improve the dense prediction tasks using multi-modal datasets [4, 15, 22, 63, 77]. A few methods [6, 43] also leverage RGB-D datasets to pre-train ViTs backbones under MAE paradigm. One such work is MultiMAE [6] which pre-trains the vision transformers using RGB, depth and segmentation masks in a multi-task MAE fashion. However, MultiMAE requires semantic segmentation labels during pre-training and access to multiple modalities during fine-tuning for downstream tasks. This method learns cross-modal relationships using a cross-attention module but under the paradigm of MAE to help reconstruct the masked patches in each modality. Another work in this line is Mask3D [43] which embeds 3D priors in ViT backbones to learn structural and geometric priors by reconstructing the masked patches in the depth input using RGB-D data. However, it doesn’t learn cross-modal relationships to capture more local context beyond 3D priors. Some prior works [20, 86, 92, 44] have also used contrastive learning for high-level scene understanding and low-level point matching [8, 91].
Motivated by this, we believe contrastive learning and Masked autoencoders can be complementary to each other as both can provide unique data representations by extracting different discriminative features from RGB-D datasets. However, MAEs alone, only focus on local spatial statistics for reconstruction. How about learning high frequency components of the data? This further motivates us to explore another learning mechanism. Recently, we have seen a remarkable success in image and video generation [41, 40, 42] using diffusion models and this success comes from the denoising technique. Inspired by this, we introduce a noise prediction loss with the masked autoencoder reconstruction objective. To do so, we add a Gaussian noise in the input and pre-train the model to predict the masked patches as well as the noise. We hypothesize that the denoising would encourage the model to extract high-frequency features from the input.
The main challenge we’re addressing in this work is how to effectively combine such methods in one single pre-training framework especially for RGB-D datasets. We have seen recent works that combine instance-level contrastive learning and masked autoencoders and leverage multi-modal data in one single framework for image recognition [47], video action recognition [60], video-text [61] and audio-video [33]. However, such methods are not suitable for RGB-D datasets.
@@
Dense prediction tasks such as semantic segmentation and depth estimation require understanding beyond image-level representation. We hypothesize that the instance-level contrastive learning would not benefit the aforementioned downstream tasks when leveraging RGB-D datasets for pre-training. So, to capture useful discriminative local context features using these modalities, patch-level or pixel-level contrastive learning would be more beneficial. However, it is not easy to combine patch-level or pixel-level contrastive learning with Masked Autoencoders in a single framework because of the masking step.
@@
Based on the above challenges, we propose a new strategy based on these self-supervised frameworks to pre-train ViTs using RGB-D datasets. Our curriculum consists of two stages, 1) contrastive learning at pixel or patch-level to align RGB and depth modalities to capture local context for cross-modal relationships 2) MIM + Denoising would encourage the encoder to learn discriminative low-frequency and high-frequency features respectively. Indeed, one can change the ordering in the curriculum, but we empirically found that learning an easier task i.e., contrastive learning is much more beneficial and improves the downstream performance. Furthermore, from the figure 1, contrastive learning converges faster compared to MIM + Denoising, making it a relatively easier task in comparison. It would also facilitate the learning to next round of pre-training (MIM + Denoising) in the curriculum. Lastly, we pre-train ViT-B on ScanNet [22] and SUN RGB-D [77] with our curriculum learning based pre-training and evaluate it on multiple downstream tasks.
Figure 1: Convergence of each learning phase of our curriculum on ScanNet RGB-D dataset. (Left) InfoNCE loss converges faster than (Right) Masked Image Modeling and Denoising making it a relatively easier task.
@@ 2 Related Work
Curriculum learning.
Humans learn information effectively when there’s a structured order in place, not by randomly presenting data samples to them. For example, there is a particular curricula for students to follow in the school. Inspired by this, the term curriculum learning in ML was first proposed in [10], providing easy-to-hard training strategies for a model. This allows the model to converge faster and achieve better performance. There has been a number of works proposing different curriculum learning strategies for various computer vision tasks such as object localization [48, 75, 79], object detection [18, 74, 82, 55] and other applications [80, 72, 36, 37, 32, 67, 51].
@@
Masked Image Modeling (MIM) has received a lot of attention for pre-training vision models since the advent of Vision Transformers (ViT). The goal of MIM is to learn discriminative features for downstream tasks by reconstructing the missing patches in the input. Masked autoencoding for vision tasks is inspired by the success of methods such as BERT [23] and GPT [69] in NLP domain. Several MIM approaches have been proposed with various reconstruction objectives based on pixels [5, 26, 16, 28, 38, 87, 50], dVAE tokens [9, 93] and features such as histogram of oriented gradients (HOG) or deep neural network embeddings [84, 7]. Specifically, MAE [38] uses an asymmetric encoder-decoder architecture based on ViT to pre-train models using masked autoencoding. It first masks large portions of an image and pass the unmasked patches into encoder followed by a decoder for reconstructing the missing patches. The masking accelerates the pre-training of ViT. Contrastive learning is different from MAE in the sense that MAE learns local spatial dependencies, while contrastive approaches learn representations such as invariance to different augmentations. This steers us to hypothesize that these two SSL frameworks learn complimentary features and therefore we combine them in a curriculum manner to get the best out of each. Indeed, one can combine them in a single framework [47, 33], but our empirical experiments show poor performance of such methods on RGB-D datasets.
@@
It involves learning from multiple modalities such as images and text [2, 52, 88, 14, 46, 21, 78], video and audio [3, 49, 62, 66], video, text and audio [1], and RGB and depth images [31]. Learning can be performed either using modality-specific encoders or shared ones. In terms of MIM, MultiMAE [6] pre-trains the ViT using RGB, depth and semantic segmentation while reconstructing the missing patches in these modalities. It also uses additional modality during the fine-tuning. Mask3D [43] uses RGB and depth modalities under MIM paradigm to encode the 3D priors by reconstructing the missing patches in the depth input.
@@ 3 Method
Our approach is a simple combination of constrastive learning, masked autoencoding and denoising used in diffusion models, designed in a curriculum for pre-training vision transformers using RGB-D data. Our goal is to learn representations from different learning mechanisms that can complement each other. The contrastive learning enhances the cross-modal representations via pixel-to-pixel level or patch-level learning, while Masked autoencoding understands the local spatial correlations, while providing 3D priors as a guidance for the RGB input. Moreover, the noise prediction focuses on the high-frequency information of the input data. We first give an overview of our approach and then explain the learning objectives i.e, contrastive, masked image modeling and noise prediction. Finally, we describe the overall pre-training curriculum pipeline.
@@ 3.1 Overview
Given a batch of $n$ RGB-D pair $\{\mathbf{x}^{rgb},\mathbf{x}^{depth}\}_{i=1}^{n}$, we pre-train ViT [26] using the combination of self-supervised frameworks in a curriculum learning manner. Our curriculum learning based pre-training consists of two stages. In the first stage, we pre-train our model using an easier task which is to learn cross-modal representations to capture correspondence using constrastive learning either at pixel-to-pixel level or patch level. In the second stage, we first initialize the encoders using the weights learned in the previous stage. Then, we use multi-modal masked autoencoder to learn more fine-grained features beyond low-level image statistics to reconstruct the missing patches in the depth input. To pre-train using masked autoencoding, we first split the RGB-D input in $T=(h/p)\times(w/p)$ patches of size $p\times p$: $\mathbf{x}_{i,\text{patch}}^{rgb},\mathbf{x}_{i,\text{patch}}^{depth}\in\mathbb{R}^{T\times p\times p\times 3}$ to pass them to the encoders. We use modality-specific encoders instead of joint encoder in the second stage. We then randomly mask patches in both inputs by keeping a percentage of patches and obtain two masks $\mathbf{M}_{i}^{rgb},\mathbf{M}_{i}^{depth}$. Following MAE [38], we pass the unmasked patches to the encoders to obtain the latent representations. The representations are then concatenated with the learnable masked tokens which are essentially the placeholders for these tokens. A lightweight decoder is used to reconstruct the masked patches. Furthermore, we also add denoising to learn the high frequency information of data. Figure 2 shows the overview of our approach.
Figure 2: Overview of our curriculum learning based pre-training. In the first stage, we pre-train the ViT-B encoder using Contrastive Learning to align the RGB and the depth patches. In the second stage, we initialize the modality-specific ViT-B encoders with the stage-1 weights and pre-train them using Multi-modal Masked autoencoding and Denoising to reconstruct the masked patches in depth input and the noise in the unmasked patches respectively.
@@ 3.2 Contrastive learning
To learn the correspondence between RGB and depth modalities, we use contrastive learning to discriminate the instances at patch-level or pixel-level. There has been a few works [58, 54] using constrastive learning to learn correspondence between RGB and pointclouds. The depth values can be indirectly related to the pointclouds and projected to 3D pointclouds using camera parameters. Therefore, we can also leverage the same framework of pixel-to-point [58] to learn the correspondence between RGB and depth. We experimented with both pixel-level and patch-level contrastive learning and found no empirical differences in the downstream performance. Specifically, for a given RGB $\mathbf{x}_{i}^{rgb}$ and depth $\mathbf{x}_{i}^{depth}$, we pass them through the shared ViT encoder to get the latent representations $\mathbf{z}^{rgb}_{i},\mathbf{z}_{i}^{depth}$. We use Conv2d as patch projection layer. Following [81], we use InfoNCE loss to pre-train the encoder with a fixed temperature $\tau$:
@@ 3.3 Masked Autoencoding
The goal of masked autoencoding is to understand local spatial statistics of the input modality. To further improve the capability of our approach, we propose to pre-train the model using MAE [38] as a second learning step in our curriculum. We initialize our modality-specific encoders with stage-1 weights. Unmasked patches of two modalities are passed to the encoders to get latent embeddings. Embeddings are then concatenated with learnable masked tokens which are then passed to the lightweight decoder. We then apply depth prediction MLP head to reconstruct the missing patches. We also experiment with RGB reconstruction task but we found that empirically it doesn’t improve the downstream performance (c.f. Table 8). Following [38, 43], we normalize the target patches and compute the loss only on masked patches. Specifically, we use MSE loss which is given as:

### v2

# A Two-Stage Progressive Pre-training using Multi-Modal Contrastive Masked Autoencoders

@@ Abstract
In this paper, we propose a new progressive pre-training method for image understanding tasks which leverages RGB-D datasets. The method utilizes Multi-Modal Contrastive Masked Autoencoder and Denoising techniques. Our proposed approach consists of two stages. In the first stage, we pre-train the model using contrastive learning to learn cross-modal representations. In the second stage, we further pre-train the model using masked autoencoding and denoising/noise prediction used in diffusion models. Masked autoencoding focuses on reconstructing the missing patches in the input modality using local spatial correlations, while denoising learns high frequency components of the input data. Moreover, it incorporates global distillation in the second stage by leveraging the knowledge acquired in stage one. Our approach is scalable, robust and suitable for pre-training RGB-D datasets. Extensive experiments on multiple datasets such as ScanNet, NYUv2 and SUN RGB-D show the efficacy and superior performance of our approach. Specifically, we show an improvement of +1.3% mIoU against Mask3D on ScanNet semantic segmentation. We further demonstrate the effectiveness of our approach in low-data regime by evaluating it for semantic segmentation task against the state-of-the-art methods.
@@ 1 Introduction
Self-supervised learning (SSL) has paved the way for improving performance in various computer vision tasks such as object recognition [37, 81, 9], semantic segmentation [43, 30, 51] and depth estimation [36] by leveraging extensive unlabeled data and large models such as Vision Transformers (ViT) [26, 54]. There are two self-supervised frameworks commonly used to pre-train large vision models. The first family of SSL is Masked Image Modeling (MIM) which is inspired by the recent success of masking in the NLP domain [23]. The seminal work in MIM is Masked Autoencoders (MAE) [37] which masks out high number of patches in the input before passing them into a Transformer encoder. Then, a lightweight decoder is used to reconstruct the masked patches in the input. Such reconstruction task provides the model with understanding of the data beyond low-level image statistics. It has achieved state-of-art performance in a variety of image recognition tasks.
@@
Recently, Masked autoencoders have been extended to pre-train transformers using multi-modal data for tasks such as video-text retrieval [53, 77], image-text retrieval [66] and video-audio-text understanding [1]. However, utilizing MAEs for dense prediction tasks like semantic segmentation and depth estimation requires understanding beyond image level statistics. With the availability of RGB-D or Time-of-Flight (ToF) cameras, various approaches [6, 43, 32, 27] have been developed to improve the dense prediction tasks using multi-modal datasets [4, 14, 21, 61, 73]. A few methods [6, 43] also leverage RGB-D datasets to pre-train ViTs backbones under MAE paradigm. One such work is MultiMAE [6] which pre-trains the vision transformers using RGB, depth and segmentation masks in a multi-task MAE fashion. However, MultiMAE requires semantic segmentation labels during pre-training and access to multiple modalities during fine-tuning for downstream tasks. This method learns cross-modal relationships using a cross-attention module but under the paradigm of MAE to help reconstruct the masked patches in each modality. Another work in this line is Mask3D [43] which embeds 3D priors in ViT backbones to learn structural and geometric priors by reconstructing the masked patches in the depth input using RGB-D data. However, it doesn’t learn cross-modal relationships to capture more local context beyond 3D priors.
The second family, contrastive learning [17, 85, 18, 35] pre-trains a model by maximizing the similarity between different augmentations of an image while minimizing the similarity to the other data samples. It essentially learns discriminative representations by enforcing invariance to data augmentations. Some prior works [19, 80, 88, 44] have explored contrastive learning for high-level scene understanding and low-level point matching [8, 87].
@@
Motivated by this, we believe contrastive learning and Masked autoencoders can be complementary to each other as both can provide unique data representations by extracting different discriminative features from RGB-D datasets. We have seen recent works that combine instance-level contrastive learning and masked autoencoders and leverage multi-modal data in one single framework for image recognition [47], video action recognition [58], video-text [59] and audio-video [33]. However, such methods are not suitable for RGB-D datasets. For example, dense prediction tasks such as semantic segmentation and depth estimation require understanding beyond image-level representation. We hypothesize that the instance-level contrastive learning would not benefit the aforementioned downstream tasks when leveraging RGB-D datasets for pre-training. So, to capture useful discriminative local context features using these modalities, patch-level or pixel-level contrastive learning would be more beneficial. However, it is not easy to combine patch-level or pixel-level contrastive learning with Masked Autoencoders in a single framework because of the masking step. Recently, CoMAE [83] proposed a single model self-supervised hybrid pre-training framework for small-scale RGB-D datasets. It incorporates contrastive learning and multi-modal masked autoencoding to pre-train a single encoder to capture cross-modal representations. However, it requires both RGB and depth data during fine-tuning and evaluation, as omitting either results in a significant drop in downstream performance. It is also limited to pre-training with small-scale datasets and fails to distill features learned in stage-1, as stage-2 continues to pre-train the encoder using only the low-level pixel reconstruction task. This might not capture rich and diverse representations, and it is an easier task and can be learned quickly because of stage-1 initialization.
Additionally, recent approaches frequently struggle to capture high-frequency components of the data, which further motivates us to explore alternative learning mechanisms. Notably, recent advancements in diffusion models have achieved remarkable success in image and video generation [41, 40, 42] by leveraging denoising techniques, a strategy that has been absent in previous approaches, prompting us to investigate its potential in our context. Inspired by this, we introduce a noise prediction loss with the masked autoencoder reconstruction objective. To do so, we add a Gaussian noise in the input and pre-train the model to predict the masked patches as well as the noise. We hypothesize that the denoising would encourage the model to extract high-frequency features from the input.
@@
Based on the above challenges, we propose a new strategy based on these self-supervised frameworks to pre-train ViTs using RGB-D datasets. Our pre-training strategy consists of two stages, 1) contrastive learning at patch-level to align RGB and depth modalities to capture local context for cross-modal relationships 2) MIM + Denoising would encourage the encoders to learn discriminative low-frequency and high-frequency features respectively. Besides, we capitalize on the feature distillation by distilling the global embeddings from the first stage, improving upon MIM + Denoising.
Lastly, we pre-train ViTs on ImageNet [22], ScanNet [21] and SUN RGB-D [73] with our progressive pre-training and evaluate it on multiple downstream tasks.
In summary, our main contributions are:
  • We present a new strategy for pre-training ViTs using RGB-D dataset. Our approach provides different flavours of self-supervised frameworks i.e, contrastive learning, masked autoencoding and feature distillation.
  • Inspired by the success of denoising in the recent diffusion models, we also propose to add a noise prediction objective to learn different discriminative features.
  • We empirically demonstrate the effectivness of our approach on multiple datasets such as ScanNet [21], NYUv2 [61] and SUN RGB-D [73] for downstream tasks like semantic segmentation and depth estimation.
  • Finally, we show the data-efficient nature of our pre-training approach by evaluating it under limited labeled data scenarios.
@@ 2 Related Work
Masked Image Modeling (MIM) has received a lot of attention for pre-training vision models since the advent of Vision Transformers (ViT). The goal of MIM is to learn discriminative features for downstream tasks by reconstructing the missing patches in the input. Masked autoencoding for vision tasks is inspired by the success of methods such as BERT [23] and GPT [67] in NLP domain. Several MIM approaches have been proposed with various reconstruction objectives based on pixels [5, 26, 16, 28, 37, 81, 49], dVAE tokens [9, 90] and features such as histogram of oriented gradients (HOG) or deep neural network embeddings [78, 7]. Specifically, MAE [37] uses an asymmetric encoder-decoder architecture based on ViT to pre-train models using masked autoencoding. It first masks large portions of an image and pass the unmasked patches into encoder followed by a decoder for reconstructing the missing patches. The masking accelerates the pre-training of ViT. Contrastive learning is different from MAE in the sense that MAE learns local spatial dependencies, while contrastive approaches learn representations such as invariance to different augmentations. This steers us to hypothesize that these two SSL frameworks learn complimentary features and therefore we combine them in a multi-stage manner to get the best out of each. Indeed, one can combine them in a single framework [47, 33], but our empirical experiments show poor performance of such methods on RGB-D datasets.
@@
It involves learning from multiple modalities such as images and text [2, 50, 82, 13, 46, 20, 74], video and audio [3, 48, 60, 64], video, text and audio [1], and RGB and depth images [32]. Learning can be performed either using modality-specific encoders or shared ones. In terms of MIM, MultiMAE [6] pre-trains the ViT using RGB, depth and semantic segmentation while reconstructing the missing patches in these modalities. It also uses additional modality during the fine-tuning. Mask3D [43] uses RGB and depth modalities under MIM paradigm to encode the 3D priors by reconstructing the missing patches in the depth input. CoMAE [83] proposes a hybrid pre-training framework consists of contrastive and masked autoencoding for small-scale RGB-D datasets.
@@
Knowledge Distillation.
Knowledge distillation [34, 39, 65] aims to transfer knowledge from one model mainly called teacher to the second model called student. It treats the output of the teacher model as the target prediction for the student model. Some approaches [39, 89, 15] use logits of the teacher model to distill knowledge while some methods [76, 84, 55, 70] leverage the intermediate features for distillation. In this paper, we present the first attempt to incorporate feature distillation for RGB-D pre-training.
@@ 3 Method
Our approach is a simple combination of constrastive learning, masked autoencoding, denoising used in diffusion models and feature distillation, designed in a two-stage pre-training for vision transformers using RGB-D data. Our goal is to learn representations from different learning mechanisms that can complement each other. We first give an overview of our approach and then explain the learning objectives i.e, contrastive, masked image modeling, noise prediction and distillation. Finally, we describe the overall progressive pre-training pipeline.
@@ 3.1 Overview
Given a batch of $n$ RGB-D pair $\{\mathbf{x}^{rgb},\mathbf{x}^{depth}\}_{i=1}^{n}$, we pre-train ViT [26] using the combination of self-supervised frameworks. Our pre-training consists of two stages. In the first stage, we pre-train our model to learn cross-modal representations to capture correspondence using constrastive learning at patch level. In the second stage, we first initialize the encoders using the weights learned in the previous stage. Then, we use multi-modal masked autoencoder to learn more fine-grained features beyond low-level image statistics to reconstruct the missing patches in the depth input. To pre-train using masked autoencoding, we first split the RGB-D input in $T=(h/p)\times(w/p)$ patches of size $p\times p$: $\mathbf{x}_{i,\text{patch}}^{rgb},\mathbf{x}_{i,\text{patch}}^{depth}\in\mathbb{R}^{T\times p\times p\times 3}$ to pass them to the encoders. We use modality-specific encoders in the both stages. We then randomly mask patches in both inputs by keeping a percentage of patches and obtain two masks $\mathbf{M}_{i}^{rgb},\mathbf{M}_{i}^{depth}$. Following MAE [37], we pass the unmasked patches to the encoders to obtain the latent representations. The representations are then concatenated with the learnable masked tokens which are essentially the placeholders for these tokens. A lightweight decoder is used to reconstruct the masked patches. Furthermore, we also add denoising to learn the high frequency information of data. Additionally, we utilize global distillation loss to efficiently leverage the knowledge acquired in the first stage. Figure 1 shows the overview of our approach.
Figure 1: Overview of our progressive pre-training. In the first stage, we pre-train the encoders using Contrastive learning to align the RGB and the depth patches. In the second stage, we initialize the modality-specific encoders with the stage-1 weights and pre-train them using Multi-modal Masked autoencoding and Denoising to reconstruct the masked patches in depth input and the noise in the unmasked patches respectively. Moreover, we incorporate Feature distillation to leverage the knowledge acquired in the stage-1.
@@ 3.2 Contrastive learning
To learn the correspondence between RGB and depth modalities, we use contrastive learning to discriminate the instances at patch-level. One can also apply contrastive learning on pixel-level by up-sampling the features to the input resolution. There has been a few works [56, 52] using constrastive learning to learn correspondence between RGB and pointclouds. The depth values can be indirectly related to the pointclouds and projected to 3D pointclouds using camera parameters. Therefore, we can also leverage the same framework for pixel-level of pixel-to-point [56] to learn the correspondence between RGB and depth. This is left for the future work. Specifically, for a given RGB $\mathbf{x}_{i}^{rgb}$ and depth $\mathbf{x}_{i}^{depth}$, we pass them through the ViT encoders to get the latent representations $\mathbf{z}^{rgb}_{i},\mathbf{z}_{i}^{depth}$. We use Conv2d as patch projection layer. Following [75], we use InfoNCE loss to pre-train the encoder with a fixed temperature $\tau$:
@@ 3.3 Masked Autoencoding
The goal of masked autoencoding is to understand local spatial statistics of the input modality. To further improve the capability of our approach, we propose to pre-train the model using MAE [37] as a second learning step. We initialize our modality-specific encoders with the stage-1 weights. Unmasked patches of two modalities are passed to the encoders to get latent embeddings. Embeddings are then concatenated with learnable masked tokens which are then passed to the lightweight decoder. We then apply depth prediction MLP head to reconstruct the missing patches. We also experiment with RGB reconstruction task but we found that empirically it doesn’t improve the downstream performance (c.f. Table 9). Following [37, 43], we normalize the target patches and compute the loss only on masked patches. Specifically, we use MSE loss which is given as: